# Efficient Decoding Methods for Language Models on Encrypted Data

### v1

@@ Appendix C Previous, Comparison-Based Methods for argmax over HE
Figure˜5 visualizes the two canonical comparison-based approaches for encrypted $\operatorname{argmax}$: the tournament method (Figure˜5(a)) and the league method (Figure˜5(b)). Both rely on repeated encrypted comparisons implemented via polynomial approximations of $\textsc{sign}(\cdot)$.
@@ Appendix D Argmax using SIMD
CKKS enables SIMD-style packing, where one ciphertext encrypts a vector $(v_{1},\ldots,v_{s})$. Operations are applied elementwise, and encrypted vectors can be rotated by $\ell<s$ positions. This allows for patterns such as $\operatorname{RotAndSum}$ Adir et al. (2024) - which aggregates values with $\log s$ rotations and additions.
@@
Figure˜6 demonstrates how to leverage SIMD to execute $n$ $\max(a,b)$ operations in parallel. Step (1) shows the encrypted input; in Step (2), we compute an indicator vector where $M[i]=1$ if $A[i]>A[i+\frac{n}{2}]$, and 0 otherwise. Subsequently, Step (3) selects (masks) only the elements that match the indicator vector, which returns the desired results. Consider two cases: $n>s$ (Figure˜6(a)) and $n=s$ (Figure˜6(b)) (The case $n<s$ behaves similarly to $n=s$). Here, $s$ is the maximal number of elements that a single ciphertext can encrypt. Figure˜6(a) illustrates the case where $n=16$ elements are encrypted in two ciphertexts, each containing a vector of 8 elements. In contrast, Figure˜6(b) shows the case $n=s$, which requires only one ciphertext. In the first case, no rotation is needed, as we can directly subtract one ciphertext from the other using element-wise subtraction. However, this is not the case in Figure˜6(b), where we must homomorphically rotate $A$ by $n/2$ positions. As a result, the output is duplicated, and we only need to keep the top $n/2$ elements. It appears that many comparisons can be saved by using SIMD. Consider the case where $n=s$ for the league method (Figure˜5(b)) that requires only $n$ homomorphic rotations and $n$ parallel sign evaluations, with one additional final parallel sign evaluation. The complexity is thus reduced from $n^{2}$ to $n$ sign operations, which, on hardware with sufficient parallel threads, may result in faster execution than the sequential tournament algorithm. When $n>s$ it is worth combining the methods as was done in Iliashenko and Zucca (2021). Starting with a tournament and when the number of elements is reduced beyond some threshold continue with the league method.
@@ Appendix E Convergence of $\operatorname{CutMax}$
This section proves convergence of $\operatorname{CutMax}$ (Algorithm˜1) in the notation of the main text. Throughout, fix $n\geq 2$, an odd integer $p\in\mathbb{Z}_{+}$ with $p\geq 3$, and a scaling constant $c\in\mathbb{R}_{+}$. Let ${Y}^{(t)}=({Y}^{(t)}_{1},\ldots,{Y}^{(t)}_{n})$ denote the state at iteration $t$, with mean and variance
$\mu^{(t)}\;:=\;\frac{1}{n}\sum_{i=1}^{n}{Y}^{(t)}_{i},\qquad(s^{(t)})^{2}\;:=\;\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}{Y}^{(t)}_{i}-\mu^{(t)}\Bigr{)}^{2},$
and standardized residuals
$r^{(t)}_{i}\;:=\;\frac{{Y}^{(t)}_{i}-\mu^{(t)}}{s^{(t)}}\,,\qquad i\in[n].$
One $\operatorname{CutMax}$ inner update (Lines 3–6 of Algorithm˜1) is
$Z^{(t)}_{i}\;:=\;1+\frac{r^{(t)}_{i}}{c},\qquad Y^{(t+1)}_{i}\;:=\;\Bigl{(}Z^{(t)}_{i}\Bigr{)}^{p},$
and the final normalization (Line 7) forms
$Z_{i}\;=\;\frac{Y^{(T)}_{i}}{\sum_{j=1}^{n}Y^{(T)}_{j}}\,$.
Let $i^{\star}:=\operatorname{argmax}_{i}Y^{(1)}_{i}$ be the unique maximizer index of the input (ties can be broken arbitrarily). The proof proceeds in four steps.
Step 1: standardization identities and sharp bounds.
For every $t$,
$\sum_{i=1}^{n}r^{(t)}_{i}=0,\qquad\frac{1}{n}\sum_{i=1}^{n}\bigl{(}r^{(t)}_{i}\bigr{)}^{2}=1.$ (7)
Moreover, the following uniform $\ell_{\infty}$ bound holds:
$\max_{i\in[n]}\bigl{|}r^{(t)}_{i}\bigr{|}\;\leq\;\sqrt{\,n-1\,}.$ (8)
Justification. The identities in (7) follow from the definitions of $\mu^{(t)}$ and $s^{(t)}$. For (8), fix $k$ and note $\sum_{i\neq k}r^{(t)}_{i}=-r^{(t)}_{k}$ by (7). By Cauchy–Schwarz, $|r^{(t)}_{k}|=\bigl{|}\sum_{i\neq k}r^{(t)}_{i}\bigr{|}\leq\sqrt{n-1}\cdot\sqrt{\sum_{i\neq k}(r^{(t)}_{i})^{2}}=\sqrt{n-1}\cdot\sqrt{n-(r^{(t)}_{k})^{2}}$, which implies $|r^{(t)}_{k}|\leq\sqrt{n-1}$.
Step 2: positivity and order preservation.
Assume
$c\;>\;\sqrt{\,n-1\,}.$ (9)
Then for every $t$ and every $i$,
$Z^{(t)}_{i}\;=\;1+\frac{r^{(t)}_{i}}{c}\;\in\;\Bigl{(}1-\frac{\sqrt{n-1}}{c},\,1+\frac{\sqrt{n-1}}{c}\Bigr{)}\;\subset\;(0,2),$
by (8). Hence $Z^{(t)}_{i}>0$ and $Y^{(t+1)}_{i}=(Z^{(t)}_{i})^{p}>0$. The maps $x\mapsto(x-\mu^{(t)})/s^{(t)}$, $u\mapsto 1+u/c$, and $v\mapsto v^{p}$ (odd $p$) are strictly increasing, and normalization by a positive scalar preserves order. Therefore,
$\operatorname{argmax}_{i}Y^{(t)}_{i}\;=\;\operatorname{argmax}_{i}Y^{(t+1)}_{i}\;=\;i^{\star}\qquad\text{for all }t\geq 1.$ (10)
Step 3: two-level manifold and its image in closed form.
Residuals are invariant under positive scaling: for any $\alpha\in\mathbb{R}_{+}$,
$\frac{\alpha{Y}^{(t)}_{i}-\alpha\mu^{(t)}}{\alpha s^{(t)}}\;=\;r^{(t)}_{i}$.
Define the normalized state $\widehat{Y}^{(t)}:=Y^{(t)}/\sum_{j}Y^{(t)}_{j}$ (so $\sum_{i}\widehat{Y}^{(t)}_{i}=1$). Let
$\mathcal{M}\;:=\;\Bigl{\{}\,y\in(0,1)^{n}:\ \sum_{i}y_{i}=1,\ y_{i^{\star}}=s,\ y_{j}=\tfrac{1-s}{n-1}\ \forall j\neq i^{\star},\ s\in(1/n,1)\Bigr{\}}$
be the two-level manifold (top coordinate $s$, non-top coordinates equal). On $\mathcal{M}$ the standardized residuals are explicit and independent of $s$:
$r_{i^{\star}}=\sqrt{n-1},\qquad r_{j}=-\frac{1}{\sqrt{n-1}}\ \ (j\neq i^{\star}),$ (11)
as the system a+(n1)b=0a+(n-1)b=0 and $a^{2}+(n-1)b^{2}=n$ has the unique solution $a=\sqrt{n-1}$ and $b=-1/\sqrt{n-1}$. Thus, for any $y\in\mathcal{M}$,
$Z_{i^{\star}}=1+\frac{\sqrt{n-1}}{c},\qquad Z_{j}=1-\frac{1}{c\sqrt{\,n-1\,}}\ \ (j\neq i^{\star}),$
and after the power and normalization (one $\operatorname{CutMax}$ inner update), the image is the same vector for all $s$:
$y^{+}\;=\;\Bigl{(}s^{\star},\ \underbrace{\tfrac{1-s^{\star}}{n-1},\ldots,\tfrac{1-s^{\star}}{n-1}}_{n-1\text{ times}}\Bigr{)},\qquad s^{\star}\;=\;\frac{\Bigl{(}1+\frac{\sqrt{n-1}}{c}\Bigr{)}^{p}}{\Bigl{(}1+\frac{\sqrt{n-1}}{c}\Bigr{)}^{p}+(n-1)\Bigl{(}1-\frac{1}{c\sqrt{n-1}}\Bigr{)}^{p}}.$ (12)
Consequently, $\mathcal{M}$ is mapped in one step to the single two-level vector (12).
Step 4: global convergence to the two–level fixed point.
We show that any trajectory with a unique maximizer converges to the two–level vector in (12).
Set–up and notation. Work with the normalized state $y$ (i.e., $\sum_{i}y_{i}=1$); this does not affect residuals or the identity of the maximizer. Let $F$ denote “one $\operatorname{CutMax}$ inner update (Lines 3–6) followed by normalization”. Fix the (unique) maximizer index $i^{\star}=\operatorname{argmax}_{i}y_{i}$. Write residuals $a:=r_{i^{\star}}(y)$ and $b_{j}:=r_{j}(y)$ for $j\neq i^{\star}$; then by (7), $\overline{b}:=\frac{1}{n-1}\sum_{j\neq i^{\star}}b_{j}=-\frac{a}{n-1}$ and all residuals lie in $\sqrt{n-1},\,\sqrt{n-1}]$ by (8). Let
$\phi(u):=\Bigl{(}1+\frac{u}{c}\Bigr{)}^{p},\qquad g_{j}:=\phi(b_{j}),\quad A:=\phi(a),\quad S:=\sum_{j\neq i^{\star}}g_{j},\quad D:=A+S.$
One $\operatorname{CutMax}$ step and normalization yield
$F(y)_{i^{\star}}=\frac{A}{D},\qquad F(y)_{j}=\frac{g_{j}}{D}\ \ (j\neq i^{\star}).$
A Lyapunov functional on residuals. Measure the dispersion of the non–tops in residual space by
$\mathcal{U}(y)\;:=\;\sum_{j\neq i^{\star}}\bigl{(}b_{j}-\overline{b}\bigr{)}^{2}\,.$
Then $\mathcal{U}(y)\geq 0$, with equality iff the non–top residuals are all equal (equivalently, $y\in\mathcal{M}$).
Bounding the non–top deviations after one step. By the mean value theorem, for each $j\neq i^{\star}$ there exists $\xi_{j}$ between $b_{j}$ and $\overline{b}$ such that
$g_{j}-\overline{g}\;=\;\phi(b_{j})-\phi(\overline{b})\;=\;\phi^{\prime}(\xi_{j})\,(b_{j}-\overline{b}),\qquad\overline{g}:=\frac{1}{n-1}\sum_{j\neq i^{\star}}g_{j}.$
Since $u\in[-\sqrt{n-1},\sqrt{n-1}]$ and $c>\sqrt{n-1}$ (assumption (9)),
$\phi^{\prime}(u)\;=\;\frac{p}{c}\Bigl{(}1+\frac{u}{c}\Bigr{)}^{p-1}\in\Bigl{[}m,\,M\Bigr{]},\quad m:=\frac{p}{c}\Bigl{(}1-\frac{\sqrt{n-1}}{c}\Bigr{)}^{p-1},\quad M:=\frac{p}{c}\Bigl{(}1+\frac{\sqrt{n-1}}{c}\Bigr{)}^{p-1}.$
Hence
$\sum_{j\neq i^{\star}}\!(g_{j}-\overline{g})^{2}\;=\;\sum_{j\neq i^{\star}}\!\bigl{(}\phi^{\prime}(\xi_{j})\bigr{)}^{2}\,(b_{j}-\overline{b})^{2}\;\leq\;M^{2}\,\sum_{j\neq i^{\star}}\!(b_{j}-\overline{b})^{2}\;=\;M^{2}\,\mathcal{U}(y).$
After normalization by the positive scalar $D$, the non–top deviations from their mean scale by $1/D$:
$\sum_{j\neq i^{\star}}\!\Bigl{(}F(y)_{j}-\frac{1-F(y)_{i^{\star}}}{n-1}\Bigr{)}^{2}\;=\;\frac{1}{D^{2}}\sum_{j\neq i^{\star}}\!(g_{j}-\overline{g})^{2}\;\leq\;\frac{M^{2}}{D^{2}}\,\mathcal{U}(y).$
Thus
$\mathcal{U}\bigl{(}F(y)\bigr{)}\ \leq\ \frac{M^{2}}{D^{2}}\,\mathcal{U}(y).$ (13)
A uniform lower bound on the normalizer. For every coordinate $u\in[-\sqrt{n-1},\sqrt{n-1}]$ we have $\phi(u)\geq\bigl{(}1-\frac{\sqrt{n-1}}{c}\bigr{)}^{\!p}$, hence
$D\;=\;A+S\;\geq\;n\Bigl{(}1-\frac{\sqrt{n-1}}{c}\Bigr{)}^{\!p}.$
Combining this with $M$, (13) gives the explicit contraction
$\mathcal{U}\!\bigl{(}F(y)\bigr{)}\ \leq\ \kappa\,\mathcal{U}(y),\qquad\kappa:=\frac{\bigl{(}\tfrac{p}{c}\bigr{)}^{2}\bigl{(}1+\tfrac{\sqrt{n-1}}{c}\bigr{)}^{2(p-1)}}{\,n^{2}\bigl{(}1-\tfrac{\sqrt{n-1}}{c}\bigr{)}^{2p}}\,.$ (14)
For the $(n,p,c)$ used in practice we have $\kappa<1$; in particular, $\kappa<1$ holds whenever $c$ is not arbitrarily close to $\sqrt{n-1}$ (for fixed $n,p$).
@@
Assume $c>\sqrt{n-1}$ and the input has a unique maximizer $i^{\star}$. Then the normalized $\operatorname{CutMax}$ iterates converge to the two–level vector in (12). Moreover, the non–top residual dispersion satisfies the geometric contraction
$\mathcal{U}\bigl{(}\widehat{Y}^{(t+1)}\bigr{)}\ \leq\ \kappa\,\mathcal{U}\bigl{(}\widehat{Y}^{(t)}\bigr{)}$
with $\kappa$ given in (14).
@@
Order preservation is Step 2; the two–level image is Step 3. In Step 4 we proved $\mathcal{U}(F(y))\leq\frac{M^{2}}{D^{2}}\,\mathcal{U}(y)$ with $M=\frac{p}{c}\bigl{(}1+\frac{\sqrt{n-1}}{c}\bigr{)}^{p-1}$ and $D\geq n\bigl{(}1-\frac{\sqrt{n-1}}{c}\bigr{)}^{p}$, which yields $\mathcal{U}(\widehat{Y}^{(t+1)})\leq\kappa\,\mathcal{U}(\widehat{Y}^{(t)})$ with $\kappa$ as in (14). Hence $\mathcal{U}(\widehat{Y}^{(t)})\!\to\!0$, so the iterates approach the two–level manifold, and by Step 3 the very next update maps any two–level point to (12). ∎
Appendix F Anecdotal Sampling Demonstration
Figure 7: Comparison of Sampling Methods on GPT-2 Last-Token Decoding: The top strip shows the complete 50,257-dimensional logit vector (only its outline is visible at this scale), and the second strip zooms in on the five largest logits. The three lower strips plot, for 1,000 repeated draws, histograms of the 30 most frequently chosen token indices produced by four different samplers: standard $\operatorname{Softmax}$ at temperature 1, the classical Gumbel-Max trick, inverse-transform sampling on the normalized exponentials (“Exp-inverse”), and our one-shot Beta-cut nucleus sampler, which perturbs each logit with i.i.d. $G_{i}\!\sim\!\mathrm{Beta}(\alpha,1)$ noise using $\alpha=0.486$ computed from ˜3.4 for $(p,q)=(0.9,0.95)$. Whereas the first three methods still allocate non-negligible mass to tail tokens, the Beta-cut histogram is sharply truncated, confirming that all probability remains inside the top-$p$ nucleus exactly as required by the HE-friendly design of Sect.˜$3.$

### v2

@@ Appendix C Previous, Comparison-Based Methods for argmax over HE
Figure˜6 visualizes the two canonical comparison-based approaches for encrypted $\operatorname{argmax}$: the tournament method (Figure˜6(a)) and the league method (Figure˜6(b)). Both rely on repeated encrypted comparisons implemented via polynomial approximations of $\textsc{sign}(\cdot)$.
@@ Appendix D Argmax using SIMD
CKKS enables SIMD-style packing, where one ciphertext encrypts a vector $(v_{1},\ldots,v_{s})$. Operations are applied elementwise, and encrypted vectors can be rotated by $\ell<s$ positions. This allows for patterns such as $\operatorname{RotAndSum}$ he4ds - which aggregates values with $\log s$ rotations and additions.
@@
Figure˜7 demonstrates how to leverage SIMD to execute $n$ $\max(a,b)$ operations in parallel. Step (1) shows the encrypted input; in Step (2), we compute an indicator vector where $M[i]=1$ if $A[i]>A[i+\frac{n}{2}]$, and 0 otherwise. Subsequently, Step (3) selects (masks) only the elements that match the indicator vector, which returns the desired results. Consider two cases: $n>s$ (Figure˜7(a)) and $n=s$ (Figure˜7(b)) (The case $n<s$ behaves similarly to $n=s$). Here, $s$ is the maximal number of elements that a single ciphertext can encrypt. Figure˜7(a) illustrates the case where $n=16$ elements are encrypted in two ciphertexts, each containing a vector of 8 elements. In contrast, Figure˜7(b) shows the case $n=s$, which requires only one ciphertext. In the first case, no rotation is needed, as we can directly subtract one ciphertext from the other using element-wise subtraction. However, this is not the case in Figure˜7(b), where we must homomorphically rotate $A$ by $n/2$ positions. As a result, the output is duplicated, and we only need to keep the top $n/2$ elements. It appears that many comparisons can be saved by using SIMD. Consider the case where $n=s$ for the league method (Figure˜6(b)) that requires only $n$ homomorphic rotations and $n$ parallel sign evaluations, with one additional final parallel sign evaluation. The complexity is thus reduced from $n^{2}$ to $n$ sign operations, which, on hardware with sufficient parallel threads, may result in faster execution than the sequential tournament algorithm. When $n>s$ it is worth combining the methods as was done in BGVCompArgmaxCombined. Starting with a tournament and when the number of elements is reduced beyond some threshold continue with the league method.
@@ Appendix E Convergence of $\operatorname{CutMax}$
E.1 Preliminaries and Definitions
The following definitions assume that a vector $x\in\mathbb{R}^{n}$ has a unique maximum element. We also denote the average of $x$ by $\mu_{x}=\frac{1}{n}\sum_{i}x_{i}$ and its standard deviation by $\sigma_{x}=\sqrt{\frac{1}{n}\sum_{i}(x_{i}-\mu_{x})^{2}}$.
Definition E.1 (Max elements).
Given a vector $x\in\mathbb{R}^{n}$ with a unique maximal element, the index of that element is denoted by ${i^{*}_{x}}=\operatorname{argmax}_{i}x_{i}$ and the index of the second-largest element (excluding the maximum) is denoted by $i^{**}_{x}=\operatorname{argmax}_{j\neq i^{*}_{x}}x_{j}$.
Definition E.2 ($\delta$-gap bound).
A vector $x\in\mathbb{R}^{n}$ with a unique maximum at index $i^{*}_{x}$ is $\delta$-gap bounded for $\delta\in\mathbb{R}_{+}$ when $x_{i^{*}_{x}}-x_{i^{**}_{x}}\geq\delta\cdot\sigma_{x}$
Definition E.3 (Gap Ratio).
For $y\in\mathbb{R}^{n}$, its gap ratio is $\Gamma(y)=\frac{y_{i^{*}_{y}}}{y_{i^{**}_{y}}}$
Definition E.4 ($c$-semi-standardization).
For $c\in\mathbb{R}_{+}$, a $c$-semi-standardization function is
$S_{c}(y):\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ (7)
$y_{i}\longmapsto\frac{y_{i}-\mu_{y}}{c\cdot\sigma_{y}}+1$ (8)
Definition E.5 ($\operatorname{CutMax}$ Iteration).
Given $c\in\mathbb{R}_{+}$ and an odd $p\in\mathbb{Z}_{+}$, the $\operatorname{CutMax}$ iteration function is
$\operatorname{CutMax}_{t}(y^{(t)}):\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ (9)
$y^{(t)}_{i}\longmapsto{\left(S_{c}(y^{(t)})\right)}^{p}$ (10)
for some iteration index $t>0$.
E.2 Key Lemmas
The following lemma shows that running a standardization on a vector $y$ preserves the ordering of elements in it. In addition, it transforms the gap condition.
Lemma E.6.
The $c$-semi-standardization function ($S_{c}$) is a strictly increasing function.
Proof.
For every $y\in\mathbb{R}^{n}$ and for every $i,j\in[n]\,,i\neq j$, because $\mu_{y},\sigma_{y},c>0$ it follows that
$y_{i}>y_{j}$ $\iff y_{i}-\mu_{y}>y_{j}-\mu_{y}$ (11)
$\iff\frac{y_{i}-\mu_{y}}{c\cdot\sigma_{y}}>\frac{y_{j}-\mu_{y}}{c\cdot\sigma_{y}}\iff S_{c}(y)_{i}>S_{c}(y)_{j}$ (12)
Lemma E.7.
Let $y\in\mathbb{R}^{n}$ be a $\delta$-gap bounded vector then $S_{c}(y)$ is a $\frac{\delta}{c}$-gap bounded vector.
Proof.
Set $z:=S_{c}(y)$, $i^{*}_{z}=i^{*}_{y}$ and $i^{**}_{z}=i^{**}_{y}$. then
$z_{i^{*}_{z}}-z_{i^{**}_{z}}=\frac{y_{i^{*}_{y}}-y_{i^{**}_{y}}}{c\cdot\sigma_{y}}\underset{\text{$y$ is $\delta$-gap bounded}}{\geq}\frac{\delta\cdot\sigma_{y}}{c\cdot\sigma_{y}}=\frac{\delta}{c}$ (13)
The next Lemma shows that applying a $c$-semi-standardization function $S_{c}$ on some vector $y$ results with negative coefficients in places where $y_{i}$ is at least $c$ standard deviations below the mean of $y$.
Lemma E.8 (Unconditional gap growth for one CutMax step).
Let ${Y}^{(t)}\in\mathbb{R}^{n}$ have a unique maximizer at $i^{*}$ and let $i^{**}\neq i^{*}$ be the runner-up index. Assume ${Y}^{(t)}$ is $\delta$-gap bounded in the sense that ${Y}^{(t)}_{i^{*}}-{Y}^{(t)}_{i^{**}}\geq\delta\,\sigma_{{Y}^{(t)}}$ (Def. E.2), where $\sigma_{{Y}^{(t)}}$ denotes the population standard deviation. Assume additionally that $\min_{i}y^{(t)}_{i}\geq\mu_{y^{(t)}}-c\sigma_{y^{(t)}}+\epsilon$ for small $\epsilon>0$, ensuring all $z_{i}>0$. Fix $c>0$ and an odd integer $p\geq 1$, and set $z=S_{c}({Y}^{(t)})=\frac{{Y}^{(t)}-\mu_{{Y}^{(t)}}}{c\,\sigma_{{Y}^{(t)}}}+1$ and $\mathbf{Y}^{(t+1)}=z^{\odot p}$ (the elementwise odd-power step of CutMax). Then the post-step gap ratio satisfies $\frac{\mathbf{Y}^{(t+1)}_{i^{*}}}{\mathbf{Y}^{(t+1)}_{i^{**}}}=\left(\frac{z_{i^{*}}}{z_{i^{**}}}\right)^{\!p}\;\;\geq\;\;\left(1+\frac{\delta}{\,c+\sqrt{n}\,}\right)^{\!p}.$ Consequently, after $T$ CutMax iterations, $\frac{y^{(T)}_{i^{*}}}{y^{(T)}_{i^{**}}}\;\geq\;\left(1+\frac{\delta}{\,c+\sqrt{n}\,}\right)^{\!pT}.$
Proof.
By Lemma E.7, $z_{i^{*}}-z_{i^{**}}\geq\delta/c.$ Therefore
$\frac{z_{i^{*}}}{z_{i^{**}}}=1+\frac{z_{i^{*}}-z_{i^{**}}}{z_{i^{**}}}\;\geq\;1+\frac{\delta}{c\,z_{i^{**}}}.$
To upper-bound $z_{i^{**}}$, write $r_{i}=\frac{{Y}^{(t)}_{i}-\mu_{{Y}^{(t)}}}{\sigma_{{Y}^{(t)}}}$. From the definition of the population standard deviation, we have $\frac{1}{n}\sum_{i}r_{i}^{2}=1\Rightarrow\sum_{i}r_{i}^{2}=n$, hence $r_{i^{**}}\leq\sqrt{n}$ and thus $z_{i^{**}}=1+\frac{r_{i^{**}}}{c}\leq 1+\frac{\sqrt{n}}{c}.$ Combining the bounds gives
$\frac{z_{i^{*}}}{z_{i^{**}}}\;\geq\;1+\frac{\delta}{c\left(1+\sqrt{n}/c\right)}=1+\frac{\delta}{\,c+\sqrt{n}\,}.$
Finally, the CutMax power step yields $\frac{\mathbf{Y}^{(t+1)}_{i^{*}}}{\mathbf{Y}^{(t+1)}_{i^{**}}}=\left(\frac{z_{i^{*}}}{z_{i^{**}}}\right)^{p}$, which gives the stated inequality. Iterating $T$ times multiplies the exponent by $T$. ∎
Lemma E.9 (Per-step gap amplification).
Let $\mathbf{y}^{(t)}\in\mathbb{R}^{n}$ satisfy the $\delta$–gap condition (Def. E.2), fix $c>0$ and odd $p\geq 1$, and set $z=S_{c}(\mathbf{y}^{(t)})$ and $\mathbf{y}^{(t+1)}=z^{\odot p}$. Then the gap ratio satisfies
$\Gamma^{(t+1)}\;=\;\Gamma^{(t)}\!\left(\frac{z_{i^{*}}}{z_{i^{**}}}\right)^{\!p}\;\geq\;\Gamma^{(t)}\!\left(1+\frac{\delta}{\,c+\sqrt{n}\,}\right)^{\!p}.$
Proof.
By Lemma E.8, $\frac{z_{i^{*}}}{z_{i^{**}}}\geq 1+\frac{\delta}{\,c+\sqrt{n}\,}$. Since $\Gamma^{(t+1)}=\Gamma^{(t)}(z_{i^{*}}/z_{i^{**}})^{p}$, the claim follows. ∎
@@ E.2 Key Lemmas
Let $\mathbf{x}\in\mathbb{R}^{n}$ satisfy the $\delta$–gap condition with $\delta>0$, and run CutMax with parameters $(p,c,T)$, $c>0$, odd $p\geq 1$. Then
$\Gamma^{(T)}\;\geq\;\Gamma^{(0)}\!\left(1+\frac{\delta}{\,c+\sqrt{n}\,}\right)^{\!pT}.$
In particular, to ensure $\Gamma^{(T)}\geq k$, it suffices to take
$T\;\geq\;\left\lceil\frac{\ln k-\ln\Gamma^{(0)}}{\,p\,\ln\!\bigl(1+\frac{\delta}{\,c+\sqrt{n}\,}\bigr)}\right\rceil\;\;\leq\;\;\left\lceil\frac{\ln k}{\,p\,\ln\!\bigl(1+\frac{\delta}{\,c+\sqrt{n}\,}\bigr)}\right\rceil,$
where the last inequality uses $\Gamma^{(0)}\geq 1$.
@@
Iterating Lemma E.9 yields $\Gamma^{(T)}\geq\Gamma^{(0)}\!\left(1+\frac{\delta}{\,c+\sqrt{n}\,}\right)^{pT}.$ Solving for $T$ gives the stated bound. ∎
Table 4: Min. iterations $T_{\min}$ for n=32,768n=32{,}768 (Mistral-7B), $p=13$, $c=5$.
$\delta$ $k=10$ $k=100$
0.5 66 132
1 34 67
2 17 34
3 12 23
5 7 14
10 4 7